Question 1 20 pts Given the information in the table, Current dividend $6.00 Growth Rate in Dividends 2.50% Required Return oplease answer
Answer #1

Using Constant Growth Model,

Stock Price in Year 1 = 6(1.025)2/(0.04 - 0.025) = $420.25

Stock Price in Year 2 = 6(1.025)3/(0.04 - 0.025) = $430.76

Capital Gain = 430.76 - 420.25

Capital Gain = $10.51